OpenStudy (anonymous):

Determine the asymptotes for the second degree equation 7x^2 +42x +6y+81=3y^2

OpenStudy (mathmale):

I'm not at all sure that the equation you've typed in has any asymptotes. It's not a rational function. Except in trigonometry, only rational functions have asymptotes. Check to ensure that you've copi

OpenStudy (mathmale):

copied this problem down correctly.

OpenStudy (anonymous):

It is a trigonometry problem. I did copy down correctly. I know it is an equation to a hyperbola and has to be manipulated. I am pretty sure it turns into to 7x^2+42x =3y^2-6y. but I am not sure where the 81 goes

OpenStudy (mathmale):

Oh, I see. In that case you mean "slant asymptotes," not vertical or horizontal ones.

OpenStudy (mathmale):

Taking the equation you've posted, re-write it as follows:\[7x^2+42x =3y^2-6y \rightarrow 7(x^2+6x)=3(y^2-2y)\]

OpenStudy (mathmale):

... as the first step towards putting this equation into "standard form" for the eq'n of a hyperbola.

OpenStudy (mathmale):

You will have to "complete the square" twice, starting with each of the expressions in each set of parentheses. Are you able to do this?

OpenStudy (anonymous):

Yes. When you complete the square you would get 7(x^2+6x+9) +3(y^2-2y+4)=81. It would then be 7(x+3)^2 + 3(y-2)^2=8. Where do you go from there?

OpenStudy (anonymous):

\[7(x^2+6x)=3(y^2-2y)\] Now move all terms to the left and side: \[7(x^2+6x)-3(y^2-2y)=0\] Complete the square for each parenthesis, \[7(x^2+6x+9)-63-3(y^2-2y+1)+3=0\] \[7(x^2+6x+9)-3(y^2-2y+1)=60\] \[7(x+3)^2-3(y-1)^2=60\] Now divide by 60 on both sides, \[\frac{(x+3)^2}{60/7}-\frac{(y-1)^2}{20}=1\] Now the equation is in the standard form

OpenStudy (mathmale):

Last, you need to determine whether this hyperbola is a vertical or a horizontal one. Once you've done that, you can identify a^2 and b^2. With acknowledgements to http://www.purplemath.com/modules/hyperbola.htm See the illustration attached:

OpenStudy (mathmale):

If you can identify your a and b correctly, the appropriate formula from the attached figure will help you identify the equations of your (slant) asymptotes.

OpenStudy (anonymous):

Well this is a horizontal asymptote since the negative sign is on y. Now you just need to find the equation of the conjugate axis, which will give you the vertical asymptote

OpenStudy (anonymous):

Sorry it is a horizontal hyperbola not a horizontal asymptote. A horizontal hyperbola has a vertical asymptote as far as I know

OpenStudy (mathmale):

Comment #1: Yes, this is a horiz. hyperbola: It opens to the left and to the right. Comment #2: The asymtopes of this hyperbola are SLANT, as I've mentioned before, not horiz. or vertical. Please see the illustration I posted a few minutes back.

OpenStudy (anonymous):

Okay. This has been so much help. I think I got it. In Navk's problem the 81 was on the other side, so I do believe when I put the 81 in the amount on the other side should be 138. The equation then is \[(x+3)^{2}/138/7 + (y-1)^{2}/46=1\]

OpenStudy (anonymous):

The asymtopes should then be \[y-1= \pm \frac{ \sqrt{46} }{ \sqrt{81/7} }(x+3)\]

OpenStudy (mathmale):

You could check the accuracy of your asymptotes by graphing them. Do that if you have lots of free time, Otherwise, I'd suggest you move on to your next homework problem.

OpenStudy (anonymous):

Okay. Perfect. Thank you!

OpenStudy (mathmale):

:)